OpenStudy (anonymous):

medal fan and testimonial asap Select the equation of the line that passes through the point (–2, 1) and has slope 3 in point-slope form. A. (x – 2) = 3(y + 1)

OpenStudy (anonymous):

@perl @AnswerMyQuestions

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

plz help me

OpenStudy (anonymous):

can u help me @AnswerMyQuestions

OpenStudy (anonymous):

I'm not sure. :(

OpenStudy (anonymous):

but your a mathlete

OpenStudy (anonymous):

please help me i wann ago to be dbut i have 2 stay up till im done

OpenStudy (anonymous):

@ilovewolf u there?

OpenStudy (ilovewolf):

Um ok \[y-y1=m(x-x1) y-1=3(x-2) y-1=3x-6 +1+1 y=3x-5\] I think thats how you figure it out I tried :/

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

@ilovewolf

OpenStudy (anonymous):

please help me

OpenStudy (ilovewolf):

I can't give you direct answers Look at the way i did it and try to figure it out :/ sorry

OpenStudy (anonymous):

pleae this is the last question before i can go to bed. plzzzzzzzzz

OpenStudy (anonymous):

@ilovewolf

OpenStudy (butterflydreamer):